### v1

# QuaLLM: An LLM-based Framework to Extract Quantitative Insights from Online Forums

Abstract.
Online discussion forums provide crucial data to understand the concerns of a wide range of real-world communities. However, the typical qualitative and quantitative methods used to analyze those data, such as thematic analysis and topic modeling, are infeasible to scale or require significant human effort to translate outputs to human readable forms. This study introduces QuaLLM, a novel LLM-based framework to analyze and extract quantitative insights from text data on online forums. The framework consists of a novel prompting methodology and evaluation strategy. We applied this framework to analyze over one million comments from two Reddit’s rideshare worker communities, marking the largest study of its type. We uncover significant worker concerns regarding AI and algorithmic platform decisions, responding to regulatory calls about worker insights. In short, our work sets a new precedent for AI-assisted quantitative data analysis to surface concerns from online forums.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; May 12, 2024; Honolulu, HI, USAisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Figure 1. Overview of QuaLLM: an LLM-Based Framework to Extract Quantitative Insights from Online Forums
Text data from online forums, such as Reddit, has been used to study issues affecting diverse communities (Proferes et al., 2021). Researchers have leveraged this data across various domains, including healthcare (Zamanifard and Robb, 2023; Xu et al., 2023; Ho et al., 2017; Paxman, 2021), political discourse (Papakyriakopoulos et al., 2023; De Francisci Morales et al., 2021; Marchal, 2020), education (Madsen and Madsen, 2022; Park et al., 2018), and gig work111a market characterized by short-term and freelance work like food delivery and ridesharing (Rosenblat and Stark, 2016; Sannon et al., 2022; Yao et al., 2021; Watkins, 2022; Ma et al., 2018), to highlight critical concerns and trends.
Studies analyzing this text data for social phenomena often use interpretive qualitative methods(Gichuru, 2017) like coding (Saldaña, 2021) and thematic analysis (Braun and Clarke, 2012), to extract themes. Although detailed, these methods demand considerable human effort and interpretation. Such detailed interpretations may not always be necessary for identifying broad trends across large online communities. Alternatively, positivist quantitative (Park et al., 2020) techniques like LDA (Abdul et al., 2018) and BERT derivatives for topic modeling also distill themes from text, but offer greater scalability. However, they require careful hyperparameter selection, operate on limited context length, and require significant human effort to translate outputs to human readable forms.
Recent advancements in LLMs and prompt engineering (Pham et al., 2023; Nori et al., 2023; OpenAI, 2023; Deng et al., 2023) offer an alternative for analyzing extensive text data from online forums, addressing limitations of scale and context. Specifically, LLMs have been shown to perform well on generation, classification, and ranking tasks (Zhao et al., 2023), and thus have the potential to assist researchers in text data analysis. However, the hitherto absence of a robust methodology and evaluation strategy for effectively using LLMs for online text data analysis limits its use and underscores a significant gap in the research landscape.
Our work addresses this gap through the following contributions:
  1. (i)
    We introduce QuaLLM, an LLM-based framework consisting of a novel prompting methodology and evaluation strategy for the analysis and extraction of quantitative insights from text data on online forums.
  2. (ii)
    We apply our framework to a case study on Reddit’s rideshare communities, analyzing over one million comments—the largest study of its kind—to identify worker concerns regarding AI and algorithmic platform decisions, responding to regulatory calls (White House (2023), OSTP).
  3. (iii)
    We discuss the broader implications of using LLMs for quantitative text data analysis.
Taken together, by establishing a new precedent for AI-assisted quantitative data analysis to surface concerns from online forums, we hope our work can significantly assist quantitative researchers across various fields.
2. Related Work
LLM Prompting and Evaluation.   LLMs have been used for generation (Pham et al., 2023; Zhao et al., 2023), classification (Deng et al., 2023), and ranking (Qin et al., 2023) tasks via advances in prompt engineering (Nori et al., 2023; OpenAI, 2023). Popular prompting methodologies include: In-context learning (Brown et al., 2020), which teaches LLMs new tasks with few labeled examples, Chain of Thought (Wei et al., 2022), which breaks down complex problems through stepwise reasoning, and Ensembling, (Wang et al., 2022) which aggregates outputs of multiple model runs to obtain a consensus output. Despite these prompting advances, accurately evaluating LLM outputs is challenging, with human evaluation generally preferred (Amer, 2023). Building on prior work, we introduce a framework consisting of a novel prompting methodology and evaluation strategy for quantitative analysis of text data from online forums.
Gig Work and Reddit Analysis.   Regulatory bodies (White House (2023), OSTP) have sought public insights on the impact of AI and algorithmic management by employers (Kantor and Sundaram, 2022; Bernhardt et al., 2021), particularly in the gig economy. Gig workers actively discuss their experiences on anonymous forums, e.g., Reddit, which have become a rich data source for researchers. Previous studies on rideshare worker concerns using Reddit, highlighting issues of privacy, scams, and support systems (Rosenblat and Stark, 2016; Sannon et al., 2022; Yao et al., 2021; Watkins, 2022; Ma et al., 2018), have been small-scale, limited by manual coding222the largest dataset analyzed about 2.6K posts (Watkins, 2022) using qualitative data analysis methods. We use our LLM-based framework to examine over 1 million comments from rideshare workers, significantly broadening the scope to amplify all voices.
3. QuaLLM Framework
We introduce QuaLLM (Figure 1), a framework consisting of a prompting methodology and evaluation strategy for the analysis of text data from online forums.
3.1. LLM Prompting Methodology for Analysis of Online Forum Data
Our approach begins with the collection of datasets from specified periods across any public threaded discussion forum, such as Reddit, Mastodon, and Facebook. We aim to extract and summarize the predominant concerns of workers within these communities through a four-step prompting process: generation, classification, aggregation, and prevalence.
3.1.1. Generation Prompt.
The generation step involves the generation of concern summaries from online discussions relevant to a specific topic and linking each concern with a representative quote. The prompt is broken down into seven steps, following best prompt engineering practices from prior research and industry (OpenAI, 2023; Adams et al., 2023; Nori et al., 2023; Wang et al., 2023). The steps include:
  1. (1)
    identifying concerns
  2. (2)
    aggregating to avoid redundancy
  3. (3)
    selecting representative quotes333we argue that forcing the model to choose and output quotes is a form of Chain-of-Thought prompting (Wei et al., 2022) strategy and find that it helps prevent hallucination
  4. (4)
    assessing concern frequency and impact
  5. (5)
    formatting outputs in JSON with concern titles, descriptions, and quotes
  6. (6)
    ensuring concerns are topic-specific and derived from input data without external knowledge
  7. (7)
    reiterating to prevent redundancy and indicating when no concerns are present
To minimize API calls, we aggregate posts and comments into groups. For each group, we programmatically enhance the JSON output with the timestamp of the earliest post and assign a unique identifier, ensuring precise identification.
All these steps are necessary since direct prompting may cause hallucinations, redundancy, and inconsistent output. This prompting strategy is similar to “open coding,” where one derives new theories and concepts informed from the underlying data (Saldaña, 2021).
3.1.2. Classification Prompt.
We define a set of primary themes, typically around 4-5, based on prior research, parallel participant interviews, or other expert input, and the outputs of the generation prompt. We then prompt the LLM to classify each identified concern accordingly, with an additional “other” category for irrelevant concerns. This step mirrors “thematic analysis” (Braun and Clarke, 2012), which groups together a number of related codes, serving as a filtration mechanism.
3.1.3. Aggregation Prompt.
Aggregates concerns into secondary themes based on classifications, and identifies the most representative concerns (e.g., top 5) within each primary theme. We can manually select representative quotes for each secondary theme.
3.1.4. Prevalence Prompt.
Classifies concerns within primary themes as belonging to one of the secondary themes or an “other” category, facilitating quantitative analysis of concern prevalence by theme percentage.
3.2. LLM Outputs Evaluation Strategy
To validate the LLM outputs we propose a mix of human and computational evaluation methods based on the following metrics: factuality and completeness for generation, accuracy for classification and prevalence, and distinctness and coverage for aggregation prompts, respectively.
3.2.1. Generation Prompt:
This prompt generates new text given unstructured source text as input.
Factuality: Determines if the LLM’s output concerns reflect the source data. Evaluators respond to: “Is this candidate concern (and associated quote) output by the LLM present in the reference human generated list of transparency concerns? Answer Yes or No”. The factuality score is the proportion of accurate LLM-generated concerns, akin to precision.
Completeness: Evaluates the LLM’s ability to capture all relevant concerns, asking: “Is this reference human concern present in the candidate concern (and associated quote) output by the LLM? Answer Yes or No.” The completeness score denotes the proportion of human-identified concerns also recognized by the LLM, akin to recall.
3.2.2. Classification and Prevalence Prompts:
Both these prompts classify source text into predefined categories.
Accuracy: Assesses the correctness of LLM classifications, measuring the proportion of accurately classified samples.
3.2.3. Aggregation Prompt:
This prompt generates aggregated text given source text as input. We leverage a topic model to generate topics for the entire text and find those topics most similar to the subthemes. Human evaluation on a subset does not accurately capture the task, and instead we pivoted to a computational evaluation on the entire text using well established topic modeling techniques.
Distinctness: We measure the proportion of unique most-similar topics associated with each sub-theme across all the sub-themes. A higher value indicates a greater diversity in the subthemes.
Coverage (k): We measure the proportion of unique most-similar topics associated with each of the ‘n’ sub-themes that are also found among the most frequent ‘nk’ topics of the overall text. For example, if n=5 and k=2, we would look at the 5 sub-themes amongst the top 10 most frequent overall topics. A higher CR-k indicates that the sub-themes represent prevalent or significant themes in the text.
4. Case Study of Worker Concerns from Rideshare Subreddits
We apply our framework to rideshare subreddit data (r/uberdrivers, r/lyftdrivers) using the GPT4-Turbo model on Microsoft Azure to identify worker concerns regarding AI and platform algorithms. See Appendices A, B, and C for our prompts, findings and costs incurred, respectively.
Rideshare subreddits were chosen for their active 400K+ member community, presence of active discussions of worker concerns; which is of interest to regulators, accessible and comprehensive data, and relevance to HCI, Economics, and Communications researchers studying the harms caused to workers as a result of rideshare platforms’ algorithmic management (Zhang et al., 2022; Dubal, 2023; Rosenblat and Stark, 2016).
We now describe our dataset and summarize our findings.
4.1. Dataset and Experiments
We obtained comprehensive 2019-2022 data from r/uberdrivers and r/lyftdrivers via https://the-eye.eu/redarcs/, focusing on this period to capture recent concerns, particularly relevant due to app changes and the unavailability of post-2022 data due to Reddit API changes [^4] Reddit API restrictions enforced in Feb’23; see https://www.redditinc.com/policies/data-api-terms. After post pre-processing to remove short samples, our dataset comprised 65,377 submissions (47,1064710647,10647 , 106 from Uber, 18,2711827118,27118 , 271 from Lyft) and 1,392,77613927761,392,7761 , 392 , 776 associated comments (1,054,03010540301,054,0301 , 054 , 030 from Uber, 338,746338746338,746338 , 746 from Lyft) [^5] On Reddit, a submission is a post, and a comment is a response to a post or another comment; we concatenate and use the text from posts and all associated comments. We then grouped submissions and comments into sets of five for each LLM API call (See Appendix A for the prompts). We encountered an 11% error rate due to throttling and content issues. Finally we obtained 58,728 concerns which we categorized into four themes: Enhancing Transparency and Explainability (24,721 concerns, 42%), Predictability and Worker Agency (12,728 concerns, 22%), Better Safety and More Time (6,144 concerns, 10.5%), and Ensuring Fairness and Non-Discrimination (4,280 concerns, 7%); the remaining 18.5% were categorized as “Other”. We then aggregated these findings to highlight the top five concerns for each theme and calculated their prevalence across the dataset.
4.2. Findings
We briefly summarize our findings below. We also provide tables of the aggregated, ranked and prevalence quantified themes in Appendix B.
(i) Enhancing Transparency and Explainability: Driver concerns about opaque fare calculations, unclear incentives, and uncertain criteria for earning, surge pricing, and cancellation (Table 1)
(ii) Predictability and Worker Agency: Drivers face unpredictable earnings from fluctuating surge pricing, algorithm changes, increased competition, complex incentive qualifications, and low compensation for long pickups and waits. (Table 2)
(iii) Better Safety and More Time: Drivers face navigation issues, support access difficulties, challenges with false complaints, and low compensation for additional tasks and wait times. (Table 3)
(iv) Ensuring Fairness and Non-Discrimination: Drivers dispute their fare share and platform take rates, fairness in ride assignments, the influence of demographics on ratings, and the criteria for deactivation. (Table 4)
4.3. Evaluation
4.3.1. Generation Prompt.
We randomly sampled 125 submissions and all the 2,511 associated comments, jointly annotated by two researchers, to create a reference list of concerns. The two researchers then jointly evaluated the LLM outputs against the reference list of concerns, and resolved any differences through discussions. We obtained a factuality score of 0.55 and a completeness score of 0.78. Lower factuality was primarily due to the LLM identifying concerns part of the input context, but not directly related to AI and algorithmic decisions. Moreover, the lower factuality score was less concerning as outputs not factual (i.e., identified by the LLM but not by humans) were mostly filtered out into the “Other” category during the classification step and excluded from any further analysis.
4.3.2. Classification and Prevalence Prompts.
Two researchers jointly annotated 100 LLM-classified concerns to establish ground truth labels, assigning one label per concern. They compared LLM outputs with these labels, achieving a consensus accuracy of 0.74. The achieved accuracy is significant because misclassified samples, excluding those labeled as ”Other,” can still contribute to the analysis. This is because concerns might fit into multiple categories depending on the analytical perspective. Even if placed in a different category, these concerns remain valuable for later aggregation and prevalence prompts. Using a similar methodology, researchers obtained an accuracy of 0.82 for the prevalence prompts’ output.
4.3.3. Aggregation Prompt.
We employed BERTopic for topic modeling (see Appendix D for hyperparameter details). Given the large volume of concerns (4K-25K) per primary theme, human evaluation on a subset would not accurately capture the task’s nuances. Therefore, we adopted this well-established computational method for evaluation. We fit a topic model on the entire text (title+description) associated with each primary theme to obtain the top 5 and 10 topics. Subsequently, we identified the most similar topic for each of the 5 sub-themes. The results demonstrated a distinctness of 0.95, coverage(1) of 0.80, and coverage(2) of 1.00, indicating that the LLMs’ outputs were both distinct and well-represented among the most frequent topics.
5. Discussion of Broader Impacts
5.1. Scaling online forum analysis with LLMs
QuaLLM presents a novel framework that leverages the power of LLMs to analyze vast amounts of unstructured text data efficiently and translate to a form which resembles community survey responses. Beyond analyzing online discussions in rideshare communities on Reddit, this methodology has broad applicability across numerous domains, across any platform, where large amounts of unstructured text is prevalent and traditional positivist quantitative methods (Park et al., 2020) such as topic models through LDA, may fall short. We now discuss a few domains where our approach can aid data analysis:
(1) Healthcare Forums::
Identify discussions of medications from patient forums comparing anecdotal experiences and widespread concerns.
(2) Customer Feedback::
Filter through customer service chats and emails to identify recurring product or service issues.
(3) Policy and Regulation::
Review public responses to policy proposals, identifying key points of support and contention.
(4) Educational Feedback::
Review online forums and course feedback to extract themes related to assessment and course content.
(5) Financial Market Sentiments::
Analyze investor forums and financial news comments to gauge market sentiment towards certain stocks or economic policies.
5.2. QuaLLM’s evaluation reveals strengths and necessitates further refinement
Due to the unique nature of our pipeline approach, which combines several tasks into a cohesive workflow, unlike the most closely related work by Pham et al. (2023) where LLMs are used for the single task of topic generation, there are no established baselines or benchmarks available for direct end-to-end comparison and evaluation. Topic models might offer a comparison for the generation prompt outputs, yet they require significant human effort to translate outputs to human readable forms, unlike our method where the LLM autonomously generates topics, making baselining challenging.
This absence of precedent opens questions regarding effective baselining and benchmarking for performance evaluation. As a result, human evaluation serves as the gold standard for evaluation, ensuring the validity of our findings.
5.3. LLMs can be used to assist researchers and complement existing methods
Some researchers highlight limitations in LLMs’ understanding of complex human emotions and experiences, pointing to a lack of context-sensitivity (Bender et al., 2021; Alkaissi and McFarlane, 2023; Rudolph et al., 2023). Others (Bano et al., 2023) argue LLMs cannot match the innate human ability to perform certain tasks in inexplicable ways, described in conceptual frameworks for knowledge creation (Li and Gao, 2003; Collins, 2005). This gap raises doubts about LLMs’ capability for interpretation crucial to fields such as anthropology (Malinowski, 1929) and sociology(Weber, 1949), challenging their application within the qualitative interpretivist research paradigm (Gichuru, 2017; Glaser et al., 1968; Denzin, 2001).
On the other hand, advocates for LLM-enhanced research suggest LLMs align with human outputs and in some cases can replace human judgement (Dwivedi et al., 2023; Bano et al., 2023; Chew et al., 2023; Xiao et al., 2023; Dai et al., 2023; Tai et al., 2023; Dunivin, 2024; Pham et al., 2023). Further, they find integrating AI with human expertise can improve the scope and scale of analyses, facilitating efficient processing of vast, unstructured datasets.
Our research stands in the middle, underscores the potential of LLMs to help researchers engage with data in a way that can supply intrinsic meaning at scale, while stressing the need to recognize their limitations and the continued importance of human oversight and interpretation.
Broadly, QuaLLM can also be used in mixed methods study designs to complement data from interviews and focus groups. However, safeguards, including successive filtering and human evaluation, are necessary to ensure validity of LLM outputs, indicating a complementary shift rather than a complete replacement of human interpretation.
6. Limitations
Interpretivist vs Positivist trade-offs::
While LLMs efficiently aggregate broad trends, reflecting a positivist quantitative approach that assumes text has intrinsic meaning, individual experiences may be neglected. This contrasts with the interpretivist qualitative paradigm, which values subjective meaning derived from the interaction between the researcher and the data.
Small scale human evaluation::
We acknowledge the limited generalizability of our human evaluation due to the relatively small sample size and its restriction to specific pipeline stages.
Hyperparameter-sensitive topic modeling::
BERTopic performance is very sensitive to choice of hyperparameters and this can result in significant variation in results based on the chosen hyperparameter values.
Reproducibility concerns::
The stochasticity of LLM outputs raises concerns about reproducibility. However, this variability mirrors the inherent subjectivity of human perspectives, which can differ across time and contexts.
7. Conclusion
Our study presents a new framework using LLMs consisting of a prompting methodology and evaluation strategy for analysis of online forum data, showcased by analyzing over one million rideshare subreddit comments—the largest analysis to date. This approach highlights LLMs’ capability to efficiently process large unstructured datasets establishing a precedent for future AI-assisted positivist quantitative research. But does an increase in data equate to enhanced understanding? Echoing Agnew et al. (2024)’s sentiments, we believe further research is essential to determine the applicability and utility of AI-assisted research methods.
Acknowledgements
We acknowledge the Azure Cloud Computing Grant from CSML at Princeton which enabled access to Microsoft Azure compute resources. We thank Sayash Kapoor for insightful discussions on the conceptual and normative implications of LLMs and prompting strategies. We are grateful to Vinayshekhar Bannihatti Kumar for his valuable inputs on defining evaluation metrics. We thank Sunnie Kim for her contributions on connections to Explainable AI. We appreciate Elizabeth Anne Watkins for providing valuable critiques from an interpretivist perspective, which helped reframe our work. We thank Yuhan Liu, Beza Desta, and Kiyosu Maeda from the Princeton HCI group for their annotations. We are grateful to Hope Schroeder for their feedback on evaluations and broad discussions on applications. We thank Prof. Arvind Narayanan and Mihir Kshirsagar from Princeton CITP for their insights on the generalization of this work and its policy implications. We thank Prof. Brian Keegan for pointing us to the Reddit data source. Finally, we thank the anonymous CSCW reviewers for their valuable feedback, which helped frame the larger paper of which this is a part.
Appendix A Prompts
GPT-4 Generation Prompt GPT-4 Classification Prompt GPT-4 Aggregation and Ranking Prompt GPT-4 Prevalence Prompt
Appendix B Findings
Harm Quote % (Count)
Drivers are unsure about opaque fare calculation methods. “\$30 to the middle of nowhere and had to eat dead miles to drive back” 29 (7,202)
Issues with understanding and qualifying for incentive programs, bonuses, and quests. “Are they really expecting a thumbnail of a map with zero details of where I’d need to be to have a ride qualify for the Quests to be useful?” 20 (4,880)
Drivers express concerns over clarity in ride assignments and earnings determination by algorithms. “A driver who keeps a high AR, low cancel rate, gets mostly 5 Stars, and gives out mostly 5 Stars to pax will get better (more lucrative) rides than a driver the company’s algorithm doesn’t like.” 11 (2,765)
Confusion persists over surge pricing determination and boundaries. “I get a ton of riders telling me they are paying 40limit-from4040-40 -60 for the ride (surging rate), but when the ride ends, I get around \$6 (non-surge rate pay)” 8 (1,858)
Concerns over lack of clarity in cancellation criteria and the associated fare impact. “Then I ask about the cancellation fee and they said they can’t provide one as no option came up.” 5 (1,197)
Table 1. Most representative concerns on Reddit relevant to the need for enhancing transparency and explainability (Total = 24,721).
Harm Quote % (Count)
Drivers face uncertain and fluctuating income due to changes in surge and quest pricing. “Uber has always done that. When they offer a higher promo, they manipulate and lower surges.” 23 (2,953)
Challenges with understanding and adapting to algorithm changes that affect ride assignments and income. “I multi-app, but the algorithm has changed. If you log off from the app, to ride for another app, the algorithm puts you on the bottom of the queue after you log on again for the next ride.” 19 (2,408)
Complexity and unpredictability in qualifying for incentive programs like quests and bonuses. “Monday thru Thursday is 20 trip min & Friday thru Sunday is 40 trip min to get the bonus. Like why?” 14 (1,799)
Concerns over inadequate payment for time and distance invested in long pickups and waits. “I don’t understand what the hell is Uber thinking when giving us long pickups for short trips. NOBODY sane accepts a \$4 ride for someone half an hour away.” 11 (1,429)
A rise in driver numbers contributing to fewer available rides and lower earnings. “Now that millions of new drivers have joined overnight due to the unemployment benefits ending, rides have dried up.” 5 (658)
Table 2. Most representative concerns on Reddit relevant to the need for enhancing predictability and worker agency (Total = 12,728).
Harm Quote % (Count)
Drivers experience technical difficulties with app navigation and route mapping, leading to disruptions. “I started using the Lyft nav, then Google maps, then Waze only to go back to the Lyft nav. I went through that cycle because the Lyft nav sucks but at least you see everything going on with the app. With the others, I’ve lost out on streaks and pings because of glitches, etc.” 20 (1,208)
Challenges in contesting false passenger complaints and inadequate platform responses affecting driver status. “They claimed the fare was adjusted due to a rider complaint or something. They’re lying.” 16 (990)
Drivers express risks due to dangerous multitasking and unsafe areas suggested by algorithms “If a Pax wants to be taken to a part of town I don’t feel safe going, I feel like I should know, for my own safety and peace of mind.” 15 (909)
Difficulty reaching and obtaining help from support for various driver safety issues “Support was really quick to respond…only took 13 days.” 14 (863)
Drivers are not fairly paid for extensive wait times during stops or added tasks “The stop system is bullshit. You make less money, 100%. I just checked. To order two rides, one there, one back, \$17. To order the same ride, but using stops it’s \$13.” 14 (856)
Table 3. Most representative concerns on Reddit relevant to the need for better safety and more time (Total = 6,144).
Harm Quote % (Count)
Drivers express discontent with the percentage of fares they receive compared to what the platform charges customers. “Yeah it’s always high ride demand and they charge the passengers high prices and drivers earnings are like 25%” 24 (1,025)
Concerns over unclear or unfair deactivation decisions based on customer complaints or automated systems. “I canceled two orders for the distance and my account was deactivated for fraud.” 20 (850)
Uncertainty about how ratings are determined and their direct impact on drivers’ work opportunities. “I read that Lyft will try to match a pax with a driver that both had 5 starred each other in the past.” 9 (384)
Drivers feel platform algorithms unfairly influence ride distribution, surge pricing, and overall income. “a lady was standing next to my car and ordered a lyft and it gave her a driver that was outside of the Bonus zone and 20 minutes away.” 7 (316)
Concern over algorithms distributing rides unfairly or prioritizing certain drivers “Perceived favoritism by the app’s algorithm towards newer drivers can limit earnings for long-term drivers.” 4 (151)
Table 4. Most representative concerns on Reddit relevant to the need for ensuring fairness and non-discrimination (Total = 4,280).
Appendix C Financial Expenditure of Using LLMs
In this section, we delve into the costs incurred from utilizing the GPT4-Turbo LLM on Microsoft Azure to synthesize concerns from the Reddit data.
Cost Category Token Quantity Cost (USD)
Input Token Rate - \$0.01 per 1K tokens
Output Token Rate - \$0.03 per 1K tokens
Total Input Tokens 135.12 million \$1351.20
Total Output Tokens 10.37 million \$311.10
Total Expenditure - \$1662.30
Table 5. GPT4-Turbo Token Usage and Associated Costs
The access to the LLM and associated compute resources through Microsoft Azure was jointly funded through a partnership between our university and Microsoft. The total expenditure for the GPT4-Turbo usage was \$1,662.30. This funding enabled us to leverage the capabilities of GPT4-Turbo to successfully meet our research objectives.
Appendix D BERTopic Hyperparameters
1# Load CSV data
2df = pd.read_csv(filename)
3
4# Concatenate title and description into one text column
5df[’text’] = df[’title’] + " " + df[’description’]
6
7dim_model = UMAP(n_neighbors=15,
8 n_components=10,
9 min_dist=0.0,
10 metric=’cosine’,
11 random_state=42)
12
13# Initialize BERTopic
14topic_model = BERTopic(
15 nr_topics="auto",
16 min_topic_size=100,
17 n_gram_range=(1, 2),
18 umap_model=dim_model
19)
20
21# Fit BERTopic on the concatenated texts from the CSV
22topic_model.fit(df[’text’])
Appendix E IRR Measurements
For the Classification and Prevalence stages, which associated a label to a given text input, we calculated inter-rater reliability (IRR). Three trained annotators annotated 100 randomly chosen samples each for both stages. We calculated human-human IRR and human-LLM IRR using Fleiss’ Kappa. To calculate human-LLM IRR, we generated the human labels by taking the majority vote of the three annotators. If no clear majority existed, i.e., all three annotators chose different labels, we labeled the class as ’Other’. The results were as follows: For the Classification stage, the human-human IRR was 0.59, and the human-LLM IRR was 0.54. For the Prevalence stage, the human-human IRR was 0.63, and the human-LLM IRR was 0.61. These results suggest good agreement according to the interpretation guidelines by Fleiss et al. (2013). Crucially, the human-human and human-LLM IRR values differed by less than 0.05, demonstrating LLMs can substitute human labels.